HOT-139: Chief Scientist Report


Chief Scientist: T. GREGORY


HOT-139: Chief Scientist Report
Chief Scientist: T. Gregory

Departed:  Aug. 26, 2002 at 1000 (HST)
Returned:  Aug. 30 at 0830 (HST)
Vessel:  R/V Wecoma
Operator:  Oregon State University
Master of the Vessel: Captain Danny Arnsdorf
Chief Scientist: Thomas Gregory
Marine Technician:  Daryl Swensen


1.  SCIENTIFIC OBJECTIVES

The objective of this cruise was to continue building a collection of
hydrographic and biogeochemical data at the Hawaii Ocean Time-series
(HOT) stations.  Four stations were to be occupied during the cruise,
in the following order:

1) Station 1, referred to as Station Kahe, is located at 21° 20.6'N, 158°
16.4'W and was to be occupied on Aug. 26 for about 3 hours.

2) Station 2: ALOHA (A Long Term Oligotrophic Habitat Assessment) is
defined as a circle with a 6 nautical mile radius centered at 22° 45'N,
158°W.  This is the main HOT station and was to be occupied for 3 days
from Aug. 27 to Aug. 29.

3) Station 8, referred to as HALE-ALOHA, is the location of our deep
ocean mooring (20° 20'N, 158° 10.6'W).  The mooring is no longer
deployed.  This station was to be occupied on Aug. 29 for about 2
hours.

4) Station 6, referred to as Station Kaena, is located off Kaena Point
at 21° 50.8'N, 158° 21.8'W and was to be occupied on Aug. 29 for about
4 hours.

A single CTD cast was to be conducted at Station 1 to collect
continuous profiles of various physical and chemical parameters.  Water
samples were to be collected at discrete depths for biogeochemical
measurements.  PRR and TSRB measurements were also to be made.

Upon arrival at Station ALOHA, net tows were to be conducted followed
by the deployment of a free-drifting sediment trap array.  After
deployment, a full-depth CTD cast was to be conducted followed by CTD
casts at strict 3-hour intervals for at least 36 hours for continuous
and discrete data collection followed by another full-depth CTD cast.
The primary production array was to be deployed on Aug. 28 for 12
hours.  Plankton net tows were to be conducted near noon and midnight
on August 27 and 28 at Station ALOHA.  PRR, TSRB and AC-9/FRRf
operations were to be done around noon Aug. 27 and 28.  The final
operation at Station ALOHA was to be the recovery of the drifting
sediment trap array.

Two Argo-style floats (Steve Riser, UW) were to be deployed at the end
of operations at Station ALOHA by D. Swift.

Following Station ALOHA operations, the ship was to transit to Station
8 to conduct one 1000 m CTD cast and then transit to Station 6.  A
near-bottom CTD cast (~2500 m) was to be conducted at Station 6
including salinity samples for calibration, after which the ship was to
transit back to Snug Harbor.

The following instruments were to collect data throughout the cruise: a
shipboard ADCP, a thermosalinograph and fluorometer, and an
anemometer.



2.  	SCIENCE PERSONNEL

PO Group:
 Shimi Rii				Research Associate	UH
 Daniel Fitzgerald (Watch Leader)	Research Associate	UH
 Mark Valenciano			Electronics Technician	UH
 Fernando Santiago-Mandujano 		Research Associate	UH
 
JGOFS Group:
 Thomas Gregory (Chief Scientist)	Research Associate	UH
 Karin Björkman				Research Specialist 	UH
 Lance Fujieki				Computer Specialist	UH
 Paul Morris				Technician		UH
 Tara Clemente				Research Associate	UH
 Cecilia Sheridan			Graduate Student	UH
 Jennifer Brum				Graduate Student	UH
 Dan Sadler (Watch Leader)		Research Associate	UH
 Andrew Hansen				Graduate Student	UH
 Ricardo Letelier 			Associate Professor	OSU
 Guido Corno				Graduate Student	OSU

Ancillary Projects:
 Dana Swift				Technician		UW


3.  GENERAL SUMMARY

All operations at Stations Kahe, ALOHA, HALE ALOHA and Kaena were
conducted as planned.  However, many of the CTD casts at Station ALOHA
were initiated later than scheduled.  This was because we were required
to run the hydro winch at a slower speed than normal, sometimes as slow
as 20 m/min during downcasts, in order to maintain adequate tension on
the wire.  Since design constraints of R/V WECOMA required CTD
deployment off the stern, the pitching motions of the ship would create
slack in the wire to a greater extent than what is experienced on other
ships when the CTD is deployed off the side.  We would get behind
schedule on casts with many bottles fired and longer duration water
sampling and then make up time on casts with fewer bottles fired and
less time-consuming water sampling.  Thirteen 1000 m and two 4800 m CTD
casts were completed at Station ALOHA.  One 1000 m cast at both
Stations Kahe and HALE-ALOHA and one 2500 m cast at Station Kaena were
obtained.  Both free-floating arrays were deployed and

C. Sheridan successfully completed six plankton net tows.

R. Letelier and G. Corno successfully completed four optical package
casts.

D. Swift successfully deployed two Argo-style profiling floats at the
end of operations at Station ALOHA.

Weather conditions were favorable throughout the cruise.

The ADCP ran without interruption throughout the cruise, as well as the
fluorometer, thermosalinograph and the ship's anemometer.

We arrived back at Snug Harbor on August 30 at around 0830.  A complete
off-load took place immediately.


4.  R/V WECOMA, OFFICERS AND CREW, TECHNICAL SUPPORT

The R/V WECOMA and her crew delivered exceptional ship support for our
work. The officers and crew were most helpful and accommodating and are
to be commended for maintaining high standards.  They showed enthusiasm
and concern for our work and were very flexible in receiving changes in
our operational schedule.  The marine technician ensured that all
operations ran smoothly and helped orient us to a vessel with which we
were not familiar.


5.  DAILY REPORT OF ACTIVITIES (HST)

August 25, 2002; Loading Day

Equipment loaded on this day.  The CTD had been used on the previous
(non-HOT) WECOMA cruise so termination was not necessary.

August 26, 2002

The ship departed from Snug harbor at 1000. Fire and abandon ship
drills were conducted at 1210, followed by a short science meeting
during which the cruise schedule was reviewed and safety issues were
discussed.

We arrived at Station Kahe at 1315 and immediately conducted PRR and
TSRB operations.  A weight cast was not necessary since the CTD had
been used on the previous cruise.  A 1000 m CTD cast was begun at 1419
and the package was back on deck at 1551.  We began transit to Station
ALOHA at 1605.

August 27, 2002

We arrived at Station ALOHA on schedule and immediately performed a net
tow followed by deployment of the sediment trap array.  The deep PO
cast started at 0225 and was back on deck at 0503.  Three bottles did
not fire so the STAG carousel was installed.  This carousel performed
well the rest of the cruise.  Following the deep cast we performed the
shallow PO cast, which initiated the 36-hr CTD cast period.  We
conducted six 1000 m casts this day.

Net tows were conducted at 0030, 1029, 1331 and 2203.

The PRR and TSRB were deployed at 1245.

August 28, 2002

Seven 1000 m CTD casts were conducted this day.  The second deep cast
was begun at 2324.

Net tows were performed at 0146 and 1000.

The PRR and TSRB were deployed at 1220.  Four AC-9/FRRf casts were
conducted at 0027, 1252, 1352 and 2143.

The primary production array was deployed at 0512 and recovered at
1920.

August 29, 2002

The deep cast was recovered at 0331 after which we steamed to the
sediment trap array and prepared for a recovery at dawn.

D. Swift successfully deployed two Argo-style floats this day at 0416
and 0727.

Upon arriving at the vicinity of the sediment trap array, the ATE
sampler was deployed at 0538.

The sediment trap array was recovered at around 0700.  The array had
drifted to the northwest.  After the sediment trap array had been
recovered we steamed to Station HALE ALOHA.

We conducted a 1000 m cast at Station HALE ALOHA then transited to
Station Kaena.  A 2500 m cast was successfully performed at Station
Kaena.  This cast was recovered at 2314 at which time we began transit
to Snug Harbor.

August 30, 2002

Arrived at Snug Harbor at around 0830.  A full offload took place
immediately.


Sub component programs:

Investigator:                   Project:
-----------------               ----------
Bob Bidigare			HPLC pigments/UH
Mike Landry			zooplankton dynamics/UH
John Dore			CO2 dynamics/UH


Ancillary programs:

Investigator:                   Project:
-----------------               ----------
Charles Keeling			CO2 dynamics and intercalibration/SIO
Mark Abbott/Ricardo Letelier	optical measurements/OSU

Others:

Investigator:                   Project:
-----------------               ----------
Karin Björkman			phosphorus cycling/UH
Steve Riser			deployment of Argo-style profiling floats/UW